Station Messages
18 Jul 2018 02:56 UTC 2018199+02:56 UTC

Overview Datums Bench Marks Data Query Site Visit Daily QC Messages Attributes

170: National Park Service - Baffin Bay

27° 17' 49" N   97° 24' 17" W
NPSBB
United States National Park Service United States National Park Service

2018073+2038 (acknowledged, lloyd)
ph : 2018069+1642 (03/10/2018 16:42 - 03/10/2018 16:43) :

Lower than expected pH values starting 2018069+1642. Probe malfunction suspect. The datasonde will be exchanged during the next site visit. pH values are being automatically removed by CBI polling.

2018158+1811 (retired, dburch)
trb : 2018089+0912-2018149+1442 (03/30/2018 09:12 - 05/29/2018 14:42) :

Turbidity reading higher than expected values from 2018089+0912-2018149+1442. Sonde probe malfunction suspect.

2018073+2036 (retired, lloyd)
do,sat : 2018059+1212-2018069+1642 (02/28/2018 12:12 - 03/10/2018 16:42) :

Dissolved oxygen spiking high from 2018059+1212 to 2018069+1642. The datasonde was replaced at 2018069+1642. The suspected cause was sediment accumulation that was found in the datasonde guard during the datasonde exchange.

2017109+1457 (retired, lloyd)
ALL : 2017042+1830 (02/11/2017 18:30 - 02/11/2017 18:31) :

Data dissemination is often lagging behind several hours, likely due to low cell signal strength. A signal booster will be installed during the next site visit.

2016172+2122 (retired, lloyd)
ALL : 2016168+1530 (06/16/2016 15:30 - 06/16/2016 15:31) :

No data dissemination starting 2016168+1530 due to datasonde cable malfunction

2016137+1152 (retired, lloyd)
2016133+0442 (05/12/2016 04:42 - 05/12/2016 04:43) :

No data dissemination starting 2016133+0442

2016120+1111 (retired, lloyd)
spcon,sal : 2016111+1912 (04/20/2016 19:12 - 04/20/2016 19:13) :

Spcon/sal suspect starting 2016111+1912. Reporting higher and lower than expected values. Probe malfunction suspect.

2016101+1422 (retired, lloyd)
2016098+0500 (04/07/2016 05:00 - 04/07/2016 05:01) :

No data dissemination starting 2016098+0500

2016075+2114 (retired, lloyd)
ALL : 2016075+0230 (03/15/2016 02:30 - 03/15/2016 02:31) :

No data being disseminated starting 2016075+0230

2016071+1819 (retired, lloyd)
spcon,sal : 2016061+2112 (03/01/2016 21:12 - 03/01/2016 21:13) :

Spcon/sal suspect starting 2016061+2112. Values occasionally reading higher and lower than expected

2016041+1458 (retired, jrizzo)

no data

2016017+1533 (retired, jrizzo)

Sal data flatline

2016005+1432 (retired, blorentson)

No data

2015349+1453 (retired, hmahlke)

no data

2015343+1740 (retired, lloyd)

No data

2015316+1615 (retired, lloyd)

No data

2015287+2117 (retired, lloyd)
ALL :

no data

2015170+1444 (retired, lloyd)

Repairs pending, waiting for replacement parts from NPS

2015091+1625 (retired, lloyd)
spcon,sal : 2015088+2012 (03/29/2015 20:12 - 03/29/2015 20:13) :

SPCOND and SAL data suspect starting 2015088+2012, lower than expected values

2015091+1624 (retired, lloyd)
spcon,sal : 2015088+1642 to 2015090+1712 ( - 01/01/1970 00:00) :

SAL and SPCOND data suspect from 2015088+1642 to 2015090+1712, lower than expected values. Low water not completely covering sensors suspected cause.

2014318+1221 (retired, amendoza)

No data

2014175+1131 (retired, amendoza)

No data, last battery reading at 9.6

2014166+1330 (retired, jrizzo)
ALL : 2014155+1300 (06/04/2014 13:00 - 06/04/2014 13:01) :

No data being disseminated starting 2014155+1300. Cannot connect to IP modem

2014156+1325 (retired, lloyd)

Turbidity suspect

2014122+1133 (retired, amendoza)

No data

2014097+1447 (retired, lloyd)

Salinity @ 0

2014028+1308 (retired, amendoza)

No data

2014026+1317 (retired, jrizzo)

Battery @ 10V

2013346+1457 (retired, lloyd)

No data

2013322+1205 (retired, amendoza)

WTP reading high

2013262+2212 (retired, lloyd)

No data. Reuploaded setup file and reset IP modem. Data collection resumed

2013231+1435 (retired, lloyd)

ph and turbidity flatline, reading 0

2013231+1435 (retired, lloyd)

WTP being removed, readings high

2013142+1841 (retired, lloyd)

All parameters are flatlining.

2013130+0857 (retired, ussery)

No data

2013120+1409 (retired, lloyd)
ALL :

Station parameters temporarily removed during station upgrade

2013107+0838 (retired, ussery)

No data since 4/15/2013 at 0100. Data has resumed.

2013044+1008 (retired, ussery)

Very little data being received since 2013042+0400.

2013044+1008 (retired, ussery)
trb,ph : 2013031+1730 (01/31/2013 17:30 - 01/31/2013 17:31) :

pH and turbidity suspect starting 2013031+1700

2013010+1036 (retired, ussery)

Little data received since 2013007+1730.

2012317+1451 (retired, lloyd)

Turbidity values reporting negative from 2012297+1800 to 2012313+1630

2012298+1407 (retired, lloyd)

Little data being received.

2012293+0000 (retired, lloyd)

All Turbidity values are negative

2012139+2159 (retired, lloyd)

Salinity data flatlining. Temperature probe failure is causing erroneous readings on all other parameters.

2012117+2152 (retired, lloyd)

No data since 4/25/2012.

2012099+0148 (retired, lloyd)

No data since 4/5/2012.

2012086+1334 (retired, lloyd)

No data since 2012073.

2012058+1045 (retired, ussery)

No data since 2/23/2012 0200. Data being received.

2012052+2204 (retired, lloyd)

No data being received after instrument swap.

2012051+1426 (retired, lloyd)

Little data received since 2012046.

2012030+1647 (retired, ussery)

No data since power outage 1/27/2012. Data collected by cbidl but not sent to lighthouse. IT has corrected the problem.

2012005+1437 (retired, lloyd)

Datasonde installed on 2011354+1200 does not report turbidity. Datasonde was swapped on 2012003+1800.

2012005+1436 (retired, lloyd)

Datasonde malfunctioning from 2011356+1530 to 2012003+1730. Datasonde was swapped.

2011361+1336 (retired, ussery)

No data received since 2011358+1200 due to server issue. Server restarted.

2011272+1926 (retired, ussery)

No data since 9/28/2011+2200. IT glitch. Data has resumed.

2011243+1448 (retired, lloyd)

Sporadic data since 8/28/2011.

2011217+1122 (retired, ussery)

No data since 2011215+1500.

2011216+1059 (retired, ussery)

DO spiking low

2011216+1059 (retired, ussery)

Turbidity spiking high

2011181+1106 (retired, ussery)

No data received since 2011171+1900.

2011161+0320 (retired, lloyd)

No data since 2011157+0430.

2011141+2140 (retired, lloyd)

No data received since 2011140+0100 due to server issues

2011108+1357 (retired, lloyd)

PH dropped to 6.5.

2011060+2150 (retired, lloyd)

DO not reporting

2011055+1826 (retired, lloyd)

No data since 2011035+0630.

2011035+1616 (retired, lloyd)

Water temp., conductivity and salinity spiking

2011005+1050 (retired, ussery)

No data since 2011004+0800. Data resumed.

2011003+1453 (retired, lloyd)

Spcon and wtp are being erratic.

2009343+1422 (retired, ussery)

Sporadic data received since 2009341.

2009288+1119 (retired, ussery)

Salinity reading very low.

2009134+1127 (retired, ussery)

Salinity is reading low.

2009082+1819 (retired, ussery)

Batt voltage high. Instrument reading voltage high.

2008348+1232 (retired, jrizzo)

YSI removed for service.

2008238+1316 (retired, ussery)

Sporadic data.

2008211+2050 (retired, ussery)

No data since 2008203.